# On $L^{p}$ extremals for Fourier extension estimate to fractional surface

### v1

@@ 4.2. Space-time localization
Now, we utilize the $L^{2}$ theory to establish the $L^{p}$-based profile decomposition for frequency localized sequences.
@@
(35) $\liminf_{n\to\infty}\left(\|f_{n}\|_{L^{p}(\mathbb{R}^{d})}^{p^{*}}-\sum_{j=1}^{J}\|\phi^{j}\|_{L^{p}(\mathbb{R}^{d})}^{p^{*}}\right)\geq 0,\quad p^{*}:=\max(p,p^{\prime});$
@@
and the smallness of remainder $L^{q}$ norms
@@
Applying Corollary 3.5, up to subsequences thereexist orthogonal parameters $\{(\xi_{n}^{j},h_{n}^{j},x_{n}^{j},t_{n}^{j})\}\subset\mathbb{R}^{d}\times\mathbb{R}_{+}\times\mathbb{R}^{d}\times\mathbb{R}$, $J_{0}\in\mathbb{N}_{+}\cup\{\infty\}$, and $\{\phi^{j}\}\subset L^{2}(\mathbb{R}^{d})$ , $\phi^{j}\neq 0$ , such that for all $J\leq J_{0}$ we have the following decomposition
@@
Meanwhile, the desired conclusions (34) and (33) ensue. For the case $p=2$, the proposition has already been demonstrated, thus our focus shifts to the scenario where $p\neq 2$.
@@
$\lim_{n\to\infty}\mathcal{E}_{\alpha}\left(e^{i(x_{n}^{j},t_{n}^{j})(\cdot,|\cdot|^{\alpha})}f_{n}\right)(x,t)=\mathcal{E}_{\alpha}\phi^{j}(x,t),\quad\quad\text{a.e.}\;\;(x,t)\in\mathbb{R}^{d+1}.$
Hence the desired orthogonality (36) follows from the classical Brézis-Lieb lemma [35, Section 1.9]. Then we turn to prove (37). Choose $p_{1}\leq p_{0}$ and $q_{1}=\frac{d+\alpha}{d}p_{1}^{\prime}$ such that $p$ lies strictly between $p_{1}$ and $2$. The size and support conditions on $f_{n}$ imply that $\{f_{n}\}\subset L^{p_{1}}(\mathbb{R}^{d})$. Then by the assumption on $p_{0}$, we have
@@
Thus for the pair $(p_{1},q_{1})$, doing the same argument as above, the Bresiz-Lieb lemma similarly implies the following $L^{q_{1}}$-orthogonality conclusion
$\left\|\mathcal{E}_{\alpha}f_{n}\right\|_{L^{q_{1}}(\mathbb{R}^{d+1})}^{q_{1}}=\sum_{j=1}^{J}\left\|\mathcal{E}_{\alpha}\phi^{j}\right\|_{L^{q_{1}}(\mathbb{R}^{d+1})}^{q_{1}}+\left\|\mathcal{E}_{\alpha}\omega_{n}^{J}\right\|_{L^{q_{1}}(\mathbb{R}^{d+1})}^{q_{1}}+o_{n\to\infty}(1).$
Hence, up to subsequences $\|\mathcal{E}_{\alpha}\omega_{n}^{J}\|_{L^{q_{1}}}$ is uniformly bounded for all $n$ and $J$. Combining this fact with the estimate (24), one can use Hölder’s inequality to deduce the desired property (37).
@@
Recalling the projection operator $\pi_{n}^{j}$ defined in Lemma 2.10 as follows
@@
Next we estimate each terms. First, the property (34) implies that
@@ 4.3. The $L^{p}$ convergence
Now we prove our main result Theorem 1.1 by using the frequency localization and the $L^{p}$-based profile decomposition obtained in the previous two subsections.
@@
Then using the properties (35)-(37) and the assumption on $p_{0}$, noticing that $q>p^{*}$, for each $m$ we conclude that
@@
$\leq M_{\alpha,p}^{p^{*}}\left(\max_{1\leq j\leq J_{0}}\|\mathcal{E}_{\alpha}\phi^{m,j}\|_{L^{q}(\mathbb{R}^{d+1})}\right)^{q-p^{*}}\sum_{j=1}^{J_{0}}\|\phi^{m,j}\|_{L^{p}(\mathbb{R}^{d})}^{p^{*}}$
$\leq M_{\alpha,p}^{p^{*}}\left(\max_{1\leq j\leq J_{0}}\|\mathcal{E}_{\alpha}\phi^{m,j}\|_{L^{q}(\mathbb{R}^{d+1})}\right)^{q-p^{*}}.$
@@
$M_{\alpha,p}^{q}-o_{m\to\infty}(1)\leq M_{\alpha,p}^{p^{*}}\|\mathcal{E}_{\alpha}\phi^{m}\|_{L^{q}(\mathbb{R}^{d+1})}^{q-p^{*}}\leq M_{\alpha,p}^{q}\|\phi^{m}\|_{L^{p}(\mathbb{R}^{d})}^{q-p^{*}}\leq M_{\alpha,p}^{q}.$
@@
Then for each fixed $m$, the lower semicontinuity of norms [35, Section 2.11] implies
@@
Then we divide the proof into two cases and will show the consequence (41) below. For the first case $p\geq 2$, the uniform convexity of $L^{p}$ norms [35, Section 2.5] implies
$A^{p}(m,n)+B^{p}(m,n)$ $\leq\left(\left\|e^{i(x_{n}^{m}t_{n}^{m})(\cdot,|\cdot|^{\alpha})}f_{n}^{m}\right\|_{L^{p}}+\|\phi^{m}\|_{L^{p}}\right)^{p}+\left|\|e^{i(x_{n}^{m}t_{n}^{m})(\cdot,|\cdot|^{\alpha})}f_{n}^{m}\|_{L^{p}}-\|\phi^{m}\|_{L^{p}}\right|^{p}$
$\leq\left(1+\|\phi^{m}\|_{L^{p}}\right)^{p}+\Big{|}1-\|\phi^{m}\|_{L^{p}}\Big{|}^{p}.$
Letting $n\to\infty$ and then $m\to\infty$, this inequality and the previous inequality (40) can directly give the desired result (41) below. For the second case $21<p<$ 2, similarly the uniform convexity of $L^{p}$ norms implies
@@
which, recalling (40) and the fact $F(x)=|x+C|^{p}$ is strictly convex for all $p>1$, further gives the desired result (41) by letting $n\to\infty$ and then $m\to\infty$.
In summary, we have shown the following consequence

### v2

@@ 4.2. Space-time localization
Now, we utilize the aforementioned $L^{2}$-theory to establish the $L^{p}$-based profile decomposition for frequency-localized sequences.
@@
(37) $\liminf_{n\to\infty}\left(\|f_{n}\|_{L^{p}(\mathbb{R}^{d})}^{p^{\prime}}-\sum_{j=1}^{J}\|\phi^{j}\|_{L^{p}(\mathbb{R}^{d})}^{p^{\prime}}\right)\geq 0;$
@@
and the smallness of the remainder $L^{q}$ norms
@@
Applying Corollary 3.5, up to subsequences there exist orthogonal parameters and
$\{(\xi_{n}^{j},h_{n}^{j},x_{n}^{j},t_{n}^{j})\}\subset\mathbb{R}^{d}\times\mathbb{R}_{+}\times\mathbb{R}^{d}\times\mathbb{R},\quad J_{0}\in\mathbb{N}_{+}\cup\{\infty\},\quad\{\phi^{j}\}\subset L^{2}(\mathbb{R}^{d}),\quad\phi^{j}\neq 0,$
such that for all $J\leq J_{0}$ we have the following decomposition
@@
Meanwhile, the desired conclusions (36) and (35) ensue.
@@
(41) $\lim_{n\to\infty}\mathcal{E}_{\alpha}\left(e^{i(x_{n}^{j},t_{n}^{j})(\cdot,|\cdot|^{\alpha})}f_{n}\right)(x,t)=\mathcal{E}_{\alpha}\phi^{j}(x,t),\quad\quad\text{a.e.}\;\;(x,t)\in\mathbb{R}^{d+1}.$
For any $p_{1}\in(1,2]$ and $q_{1}=\frac{d+\alpha}{d}p_{1}^{\prime}$, the size and support conditions on $f_{n}$ imply that $\{f_{n}\}\subset L^{p_{1}}(\mathbb{R}^{d})$. Then by the boundness of $\mathcal{E}_{\alpha}$, we have
@@
Combining this with pointwise estimate (41), the classical Brézis-Lieb lemma [31, Section 1.9] implies the following $L^{q_{1}}$-orthogonality conclusion
(42) $\left\|\mathcal{E}_{\alpha}f_{n}\right\|_{L^{q_{1}}(\mathbb{R}^{d+1})}^{q_{1}}=\sum_{j=1}^{J}\left\|\mathcal{E}_{\alpha}\phi^{j}\right\|_{L^{q_{1}}(\mathbb{R}^{d+1})}^{q_{1}}+\left\|\mathcal{E}_{\alpha}\omega_{n}^{J}\right\|_{L^{q_{1}}(\mathbb{R}^{d+1})}^{q_{1}}+o_{n\to\infty}(1),$
which includes the desired orthogonality (38). Choosing $p_{1}\in(1,p)$, the $L^{q_{1}}$-orthogonality conclusion (42) also implies that up to subsequences $\|\mathcal{E}_{\alpha}\omega_{n}^{J}\|_{L^{q_{1}}}$ is uniformly bounded for all $n$ and $J$. Combining this fact with the estimate (24), one can use Hölder’s inequality to deduce the desired property (39).
@@
Recalling the projection operator $\pi_{n}^{j}$ in Lemma 2.10 defined as follows
@@
Next we estimate each term. First, the property (36) implies that
@@ 4.3. The $L^{p}$ convergence
Here we prove our main result Theorem 1.1 by using the frequency localization and the $L^{p}$-based profile decomposition obtained in the previous two subsections.
@@
Then using the properties (37)-(39) and $L^{q_{1}}$-orthogonality conclusion, noticing that $q>p^{\prime}$, for each $m$ we conclude that
@@
$\leq M_{\alpha,p}^{p^{*}}\left(\max_{1\leq j\leq J_{0}}\|\mathcal{E}_{\alpha}\phi^{m,j}\|_{L^{q}(\mathbb{R}^{d+1})}\right)^{q-p^{\prime}}\sum_{j=1}^{J_{0}}\|\phi^{m,j}\|_{L^{p}(\mathbb{R}^{d})}^{p^{\prime}}$
$\leq M_{\alpha,p}^{p^{*}}\left(\max_{1\leq j\leq J_{0}}\|\mathcal{E}_{\alpha}\phi^{m,j}\|_{L^{q}(\mathbb{R}^{d+1})}\right)^{q-p^{\prime}}.$
@@
$M_{\alpha,p}^{q}-o_{m\to\infty}(1)\leq M_{\alpha,p}^{p^{\prime}}\|\mathcal{E}_{\alpha}\phi^{m}\|_{L^{q}(\mathbb{R}^{d+1})}^{q-p^{\prime}}\leq M_{\alpha,p}^{q}\|\phi^{m}\|_{L^{p}(\mathbb{R}^{d})}^{q-p^{\prime}}\leq M_{\alpha,p}^{q}.$
@@
Then for each fixed $m$, the lower semicontinuity of norms [31, Section 2.11] implies
@@
The uniform convexity of $L^{p}$ norms [31, Section 2.5] implies
@@
which, recalling (44) and the fact $F(x)=|x+C|^{p}$ is strictly convex for all $p>1$, further gives the following result (45) by letting $n\to\infty$ and then $m\to\infty$,